OpenStudy (anonymous):

Solve log7 (x – 8) – log7 3 = log7 21

OpenStudy (anonymous):

log7 (x – 8) – log7 3 = log7 21 Condition x > 8 log7(x - 8) = log7(21) + log7(3) log7(x - 8) = log7(21*3) Anti log x - 8 = 63 x = 72

OpenStudy (anonymous):

Thank you!(: